Multi-instrument study of the solar wind-magnetosphere-ionosphere coupling

Contact: Dr. Frédéric Pitout ; fpitout@rssd.esa.int status: oral – time: 14:15 – Invited

Simultaneous convection measurements by EDI on Cluster and SuperDARN radars

Contact: Dr. Matthias Förster ; mfo@mpe.mpg.de status: oral – time: 14:45 –

Cluster Observations of ELF/VLF signals generated by modulated heating of the lower ionosphere with the HAARP HF transmitter

Contact: Prof. Umran S. Inan ; inan@nova.stanford.edu status: oral – time: 15:00 –

Local and global ionic structures inside the Ring Current region : CLUSTER/CIS and FGM and IMAGE/HENA observations

Contact: VALLAT ; vallat@cesr.fr status: oral – time: 15:15 –

Spatial/temporal coherence of dayside oxygen ion outflows observed by Cluster

Contact: M. Bouhram ; bouhram@mpe.mpg.de status: oral – time: 15:30 –

FAST–CLUSTER conjunctions above the auroral oval

Contact: MSc. Octav Marghitu ; marghitu@venus.nipne.ro status: oral – time: 15:45 –

Remote sensing with wave observations: An introduction

Contact: CANU Patrick ; Patrick.Canu@cetp.ipsl.fr status: oral – time: 16:20 – Invited

Spatial and Temporal Properties of AKR Burst Emission Derived From WBD VLBI Studies

Contact: Prof. Robert Mutel ; robert-mutel@uiowa.edu status: oral – time: 16:35 – Invited

Four spacecraft determination of wave front normal and velocity

Contact: G. Chanteur ; Gerard.Chanteur@cetp.ipsl.fr status: oral – time: 17:00 –

Multipoint analysis of the source region of storm-time chorus

Contact: Ondrej Santolik ; ondrej.santolik@mff.cuni.cz status: oral – time: 17:15 –

Multi-point measurements of the source structure and motion of discrete electromagnetic waves

Contact: Prof. Umran S. Inan ; inan@nova.stanford.edu status: oral – time: 17:30 –

Continuum radiations characteristics close and far from sources: case events observations from the four CLUSTER spacecraft

Contact: Dr Pierrette Décréau ; pdecreau@cnrs-orleans.fr status: oral – time: 17:45 –

Magnetospheric Radio Tomography Experiments using Cluster and IMAGE

Contact: Steven Cummer ; cummer@ee.duke.edu status: poster – time: 00:00 –

Multipoint observations of ionic structures in the Plasmasphere by CLUSTER - CIS and comparisons with IMAGE-EUV observations

Contact: Dr. Iannis Dandouras ; Iannis.Dandouras@cesr.fr status: poster – time: 00:00 –

Cluster multipoint observations of high altitude oxygen ion energization and outflow

Contact: Dr Hans Nilsson ; hans.nilsson@irf.se status: poster – time: 00:00 –

On the evolution of transversely heated oxygen ion distributions along the cusp/cleft field lines

Contact: Dr. Bouhram Mehdi ; bouhram@mpe.mpg.de status: poster – time: 00:00 –

Characteristics of chorus emissions observed by the CLUSTER satellites

Contact: PARROT ; mparrot@cnrs-orleans.fr status: poster – time: 00:00 –

Statistical study of equatorial noise below the local lower hybrid frequency observed by Cluster

Contact: Ondrej Santolik ; ondrej.santolik@mff.cuni.cz status: poster – time: 00:00 –

2Fpe Emissions Observed in the Earth’s Foreshock Regions by the WHISPER Experiment Onboard CLUSTER: Event Studies

Contact: Dr. TROTIGNON J. G. ; jgtrotig@cnrs-orleans.fr status: poster – time: 00:00 –